OpenStudy (anonymous):

What is the y-intercept of the graph: 3x – 9y = 18. A. 0,2 b.0,-2 is it A.

jimthompson5910 (jim_thompson5910):

Now plug in x = 0 and solve for y 3x – 9y = 18 3(0) – 9y = 18 0 - 9y = 18 -9y = 18 y = 18/(-9) y = -2 So the y-intercept is (0,-2)

jimthompson5910 (jim_thompson5910):

So you were close, just off by a sign

OpenStudy (anonymous):

okay thank you. i wasnt sure if the negative was with the 9.

jimthompson5910 (jim_thompson5910):

yes subtract 9y is the same as adding -9y

OpenStudy (anonymous):

got it. thank you

jimthompson5910 (jim_thompson5910):

yw